OpenStudy (anonymous):

Help please. The water level varies from 12 inches at low tide to 52 inches at high tide. Low tide occurs at 9:15 a.m. and high tide occurs at 3:30 p.m. What is a cosine function that models the variation in inches above and below the water level as a function of time in hours since 9:15 a.m.?

OpenStudy (anonymous):

@jim_thompson5910 @Hero @.Sam. @robtobey @hba @campbell_st

OpenStudy (campbell_st):

well the difference in the tide is 40 half of that value gives the amplitude a = 20 so you have \[f(x) = 20\cos(bx) + c\] the value of c is 32... since the centre of the has been moved up 32 units so that the minimum amplitude is 12 = ( 32 - 20) and the maximum is 52 = (32 + 20) so now you get \[f(x) = 20\cos(bx) + 32\] so if the curve takes 6 1/4 hours from low to high tides then it will take 12 1/2 hours to complete a full cycle. so now you need to adjust the period... converting 12 1/2 hours to an angle measure.

OpenStudy (anonymous):

375 degrees? I'm assuming I like envision a clock to convert that to an angle measure but I could be completely wrong

OpenStudy (campbell_st):

oops should be 12/12.5 for the value of b

OpenStudy (rajee_sam):

The period will be 12.5 So \[f(x) = -20 \cos [\frac{ 2\pi }{ 12.5 }(x-9.25)] + 32\]

OpenStudy (anonymous):

I'm confused, is that my final answer? ^

OpenStudy (rajee_sam):

the cosine curve starts at the low tide so -20 ; period is 12.5 so 2pi /12.5 and the curve is shifted 9.25 to the right so (x-9.25) and it is shifted 32 up so +32. Am I right @campbell_st

OpenStudy (rajee_sam):

I am 99 % sure that is right

OpenStudy (anonymous):

@campbell_st I'm confused, what do I plug in for the value of b? .96? (12/12.5)

OpenStudy (rajee_sam):

consider the clock has 24 segments for 24 hours. So one whole day is 2pi one full rotation. you divide that into period of 12.5 hrs. so 2pi / 12.5

OpenStudy (anonymous):

How come he has 20cos up above and you have -20cos?

OpenStudy (rajee_sam):

and x is not just x because the graph is transformed to the right meaning shifted to the right by 9.25 ( starting of the low tide)

OpenStudy (rajee_sam):

-20 meaning it starts with the low tide. The middle is 32 , the low point is at 12 which is 20 down so -20 then it goes up.

OpenStudy (anonymous):

Ok, thanks :)... one more thing, my problem asks me to show work, do you know what kind of work I can show for this problem?

OpenStudy (rajee_sam):

you can draw the graph

OpenStudy (anonymous):

I can't draw anything in my applet

OpenStudy (rajee_sam):

you can use explanation given my @campbell_st with the minor corrections I gave and show your work that way

OpenStudy (rajee_sam):

for the most part of it he is right. how he got the 20, the 32 etc. You add the -ve sign and the explanation I gave you for finding the b and also the transformation of x to x-9.25 etc. and write you answer

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

@rajee_sam I really think that it is positive 20 cos not -20.... the amplitude can't be negative

OpenStudy (rajee_sam):

amplitude is not negative the direction is negative

OpenStudy (anonymous):

y=ACos(Bx-C) A=amplitude B=coefficient of x Period=2pi/B Phase shift=C/B

OpenStudy (anonymous):

I don't see how the direction matters in the equation, it's just plug in for a, or the amplitude, which is 20

OpenStudy (rajee_sam):

 in a day the first one that starts the cos curve is the low tide which is -20 from the mean